You are considering purchasing your first home. The price is $160,000, and you have saved up a $10,000 down payment. If you can get a mortgage loan at 5.2%, what will be your monthly payment to repay this loan?
Answer #1

Using BA II plus texas instrument calculator,

PV=($10,000)

FV=$160,000

N=12(since monthly payments have to be calculated)

I/Y=5.2%/12=0.43% p.a compounded monthly

CPT-PMT=$12,162
Answer #2

Solution: Calculating Monthly Mortgage Payment


Step 1: Determine Loan Amount

  • Home price: $160,000

  • Down payment: $10,000

Loan amount=160,00010,000=150,000\text{Loan amount} = 160,000 - 10,000 = 150,000


Step 2: Define Variables

  • Loan amount (P)(P) = $150,000

  • Annual interest rate (r)(r) = 5.2% or 0.052

  • Monthly interest rate (rm)(r_m):

rm=0.05212=0.00433r_m = \frac{0.052}{12} = 0.00433

  • Loan term (n)(n):

n=30×12=360monthsn = 30 \times 12 = 360 \, \text{months}


Step 3: Use the Mortgage Payment Formula

The formula to calculate the monthly mortgage payment is:

M=P×rm×(1+rm)n(1+rm)n1M = \frac{P \times r_m \times (1 + r_m)^n}{(1 + r_m)^n - 1}

Where:

  • MM is the monthly payment

  • P=150,000P = 150,000

  • rm=0.00433r_m = 0.00433

  • n=360n = 360


Step 4: Substitute the Values and Calculate

M=150,000×0.00433×(1+0.00433)360(1+0.00433)3601M = \frac{150,000 \times 0.00433 \times (1 + 0.00433)^{360}}{(1 + 0.00433)^{360} - 1}

After performing the calculation:

M823.67M \approx 823.67


Step 5: Final Answer

The monthly payment to repay this loan will be approximately $823.67.\text{The monthly payment to repay this loan will be approximately } \mathbf{\$823.67.}


answered by: Monu Kumar Gupta
Answer #3

To determine your monthly mortgage payment for a 160,000homewitha10,000 down payment and a 5.2% interest rate, here's a breakdown:

  1. Loan Amount:
    160,000(homeprice)10,000 (down payment) = $150,000.

  2. Interest Rate and Term:

    • Annual interest rate: 5.2%

    • Monthly interest rate: 5.2%120.4333% or 0.004333 in decimal.

    • Assumed loan term: 30 years (360 months).

  3. Monthly Payment Formula:

    M=P×r(1+r)n(1+r)n1

    Where:

    • P=$150,000 (loan amount)

    • r=0.004333 (monthly interest rate)

    • n=360 (number of payments).

  4. Calculation:
    Plugging in the values, the monthly payment calculates to approximately:
    $823.08.

Final Answer:
Your monthly mortgage payment would be $823.08.


answered by: anonymous